Emireee:

What does 7(3x+8) equal?

snowflake0531:

The distributive property says a(b+c) = ab+ac So do the multiplication

snowflake0531:

\[7(3x+8) = 7(3x) + 7(8)\]

Emireee:

so the first one would be 21x?

snowflake0531:

yes 21x+ what

jake111:

Im lost

Emireee:

56?

Tyrion:

yess

snowflake0531:

yes 21x+56

Emireee:

yay thx

snowflake0531:

yw~